OpenStudy (sagewilson):

Will medal. Standard Form Indicate in standard form the equation of the line passing through the given points. G(4, 6), H(1, 5) Slope: 1/-3 Answer __ - __ __ __ - __ __ How would I solve this?

OpenStudy (jessicaloraine):

y= -1/3m+ b and to find b, you need to use point slope form I think

OpenStudy (sagewilson):

Yes the point slope form, Y - Y1 = M (X - X1) I have done that but I don't think I'm doing it right.

OpenStudy (sagewilson):

i don't know how I would go about doing this given that I need to do the distributive property 1/-3 * X, 1/-3 * 4. I know I would end up with Y - 6 = 1/-3 - ??